### v2

# Towards Safe and Robust Autonomous Vehicle Platooning: A Self-Organizing Cooperative Decision-making Framework

Abstract
In hybrid traffic environments where human-driven vehicles (HDVs) and autonomous vehicles (AVs) coexist, achieving safe and robust decision-making for AV platooning remains a complex challenge. Existing platooning systems often struggle with dynamic formation management and adaptability, especially in unpredictable, mixed-traffic conditions. To enhance autonomous vehicle platooning within these hybrid environments, this paper presents TriCoD, a twin-world safety-enhanced Data-Model-Knowledge Triple-Driven Cooperative Decision-making Framework. This framework integrates deep reinforcement learning (DRL) with model-driven approaches, enabling dynamic formation dissolution and reconfiguration through a safety-prioritized twin-world deduction mechanism. The DRL component augments traditional model-driven methods, enhancing both safety and operational efficiency, especially under emergency conditions. Additionally, an adaptive switching mechanism allows the system to seamlessly shift between data-driven and model-driven strategies based on real-time traffic demands, thereby optimizing decision-making ability and adaptability. Simulation experiments and hardware-in-the-loop tests demonstrate that the proposed framework significantly improves safety, robustness, and flexibility. A detailed account of the validation results for the model can be found in Our Website.
Index Terms:
Autonomous Vehicle Platooning; Deep Reinforcement Learning; Mixed Traffic Flows; Self-Organised Cooperative Decision-making.
I Introduction
The rapid advancements in autonomous driving and vehicle communication technologies have catalyzed significant interest in connected and autonomous vehicles (CAVs). Among these technological advances, autonomous vehicle platooning has emerged as a promising approach to enhance traffic efficiency, safety, and environmental sustainability [1, 2]. By reducing inter-vehicle gaps and aerodynamic drag, vehicle platooning holds substantial potential to decrease fuel consumption and emissions[3], addressing environmental challenges and mitigating traffic congestion[4, 5].
Current autonomous vehicle platooning systems primarily focus on two key aspects of vehicle control: Cooperative Adaptive Cruise Control (CACC) and Cooperative Lane Change Control (CLCC). CACC manages longitudinal dynamics to maintain inter-vehicle gaps and ensure string stability [6, 7], while CLCC facilitates lateral maneuvers for lane changes within the platoon. Extensive research has been dedicated to improving CACC and CLCC. Model-based approaches, such as Model Predictive Control (MPC), have been employed to optimize vehicle dynamics [8, 9, 10]. However, these approaches often struggle with computational complexity and adaptability to rapidly changing environments. Rule-based methods offer simplicity but lack flexibility [11, 12, 13], while Reinforcement Learning (RL) has recently emerged as a promising alternative capable of learning from real-time traffic data [7, 14].
Despite these contributions, CACC and CLCC face intrinsic limitations in managing dynamic interactions and formation reconfiguration within mixed traffic environments[15, 16], restricting their applicability to real-world scenarios characterized by complex and unpredictable conditions [17]. Addressing these challenges necessitates a more dynamic and adaptable approach to formation maintenance and reconfiguration.
To overcome these challenges, inspiration is drawn from formation decision-making techniques in other domains, specifically Unmanned Aerial Vehicles (UAVs). In the broader context of multi-agent systems, the concept of formation reconfiguration has been extensively explored in UAVs [18], which leverage dynamic formation reconfiguration to adapt to changing conditions, avoid obstacles, and fulfill mission requirements in controlled three-dimensional environments[19, 20]. Although UAVs operate in different environments compared to ground vehicles, certain principles, such as decentralized control and dynamic adaptability, can be applied to autonomous vehicle platoons to improve their formation decision capabilities.
Unlike UAVs, which operate in three dimensions with relatively low obstacle densities, autonomous vehicle platoons must navigate constrained road environments with frequent interactions involving human-driven vehicles (HDVs) and complex infrastructure[21]. The unpredictable behavior of HDVs, combined with road constraints, presents a significant challenge for dynamic formation reconfiguration in vehicle platoons[22]. Addressing these challenges requires a novel approach that integrates adaptability with the specific needs of ground-based traffic systems.
In response, this paper proposes TriCoD, a safety-enhanced Data-Model-Knowledge Triple-Driven Adaptive Cooperative Decision-making Framework for autonomous vehicle platooning, integrating data-driven methods with model-based control to manage platoons in complex environments. The platooning problem is modeled as a multi-agent cooperative decision-making challenge based on Trust Region Policy Optimization (TRPO) [23], with a twin-world safety enhancement model runs in parallel with real scenarios to improve safety and adaptability during learning. An adaptive switching mechanism optimizes transitions between decision strategies based on traffic conditions.
The hybrid approach is designed to enable platoon reconfiguration in complex traffic conditions. The system dynamically adjusts both longitudinal and lateral behaviors to enhance safety, efficiency, and flexibility in real time. The approach is evaluated in high-risk mixed-traffic scenarios, including human interference, incidents, and fluctuating flows, as shown in Figure 1. Extensive software-in-the-loop and hardware-in-the-loop experiments fully demonstrate that our framework exhibits higher security and robustness in different scenarios while taking efficiency into consideration.
The contributions of this paper are summarized as follows:
  • A hybrid framework for platooning decision has been developed, which integrates data-driven methods with model-based control and domain-specific knowledge, enabling platoon members to achieve adaptive decision-making in complex traffic environments.
  • A twin-world safety enhancement model integrated with RL is proposed, which significantly enhances the sampling efficiency by providing a parallel virtual environment that allows safe exploration of diverse strategies, thereby improving the overall safety and adaptability of the system.
  • The proposed framework uniquely enables the dissolution and reconfiguration of platoon formations in response to changing traffic conditions. It integrates longitudinal and lateral control, allowing the system to dynamically adjust vehicle behaviors and maintain formation integrity in complex environments.
To better present our work, the rest of this paper is organized as follows. Section II formulates the problem in our research. Section III introduces in detail the proposed TriCoD framework. Section IV presents the performance evaluation. Finally, the paper is concluded in Section V.
Figure 1: Illustration of challenging traffic conditions and our framework’s adaptive responses: formation maintenance, dissolution, and reorganization.CAVs (red) and HDVs (green) are represented with corresponding behaviors under each scenario.
II Problem statement
II-A Scenario Description
Highway environments are the most common application scenario for autonomous vehicle platooning. This study focuses on self-organizing decision-making for multi-vehicle platooning on a three-lane highway, addressing significant disturbances such as human interference, traffic accidents, and traffic flow oscillations, as illustrated in Figure 1. In these scenarios, traditional formation algorithms are no longer capable of effectively addressing the challenges posed by unpredictable disturbances, leading to reduced stability and compromised safety. Therefore, a new, adaptive algorithm is necessary to navigate these disturbances and ensure robust, stable, high-speed operation.
II-B Problem Formulation
The problem is modeled as a Partially Observable Markov Decision Process (POMDP)[24], which provides a robust framework for handling the partial observability and stochastic nature of real-world traffic scenarios. By employing DRL within the POMDP framework, both individual and collective goals can be balanced, enabling adaptive responses in dynamic environments. The system’s schematics are shown in Figure 2.
The vehicle platoon, while making decisions as a whole, employs a decentralized execution strategy. The problem is formulated as a POMDP, which can be described by the tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{O})$, where $\mathcal{S}$ is the set of system states; $\mathcal{A}$ is the action space; $\mathcal{P}$ is the state transition model; $\mathcal{R}$ is the reward function; and $\mathcal{O}$ is the observation space.
In this POMDP, the agent makes decisions based on the policy $\pi$ and the current observation space. The optimization problem is then formulated as finding the optimal policy $\pi^{*}$ that maximizes the expected sum of discounted future rewards[25]:
$Q^{\pi}(s,a)\stackrel{{\scriptstyle\text{def}}}{{=}}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]$ (1)
where $\gamma\in[0,1]$ is the discount factor that balances the importance of immediate versus future rewards. The optimization process involves updating the policy $\pi_{\theta}$ to approximate the optimal policy $\pi^{*}$, which governs how the platoon should act to achieve long-term performance goals.
II-B1 Observation Space $\mathcal{O}$
In real-world highway environments, autonomous vehicles cannot directly observe the full system state due to sensor limitations and occlusions caused by other vehicles. Formally, the observation $o_{t}$ at time $t$ is a function of the true state $s_{t}$, but is incomplete:
$o_{t}=f(s_{t}),\quad o_{t}\subset s_{t}$ (2)
In this study, the observation space is structured as a matrix with dimensions $|N_{i}|\times|F|$, where $|N_{i}|$ is the number of vehicles detected by the platoon, and $|F|$ represents the feature vector describing each detected vehicle. The feature vector includes relative positions and velocities of other vehicles, as well as an indicator variable $\delta_{i}^{platoon}$ that specifies whether the vehicle is part of the platoon:
$F_{i}=[x_{i},y_{i},v_{i}^{x},v_{i}^{y},\delta_{i}^{platoon}]$ (3)
where $x_{i}$ and $y_{i}$ represent the relative longitudinal and lateral positions of vehicle $i$, and $v_{i}^{x}$ and $v_{i}^{y}$ denote its relative longitudinal and lateral velocities. The binary variable $\delta_{i}^{platoon}$ indicates whether vehicle $i$ is part of the platoon:
$\delta_{i}^{platoon}=\begin{cases}1&\text{if vehicle }i\text{ belongs to the platoon},\\ 0&\text{otherwise}.\end{cases}$ (4)
Additionally, the observation is restricted to vehicles within a sensor detection range, $D_{vision}$, ahead or behind the observing vehicle.
Figure 2: Illustration of the system and the simulation setup. All high-level actions are verified through a security mask before being processed by the bottom-level controller, which will be finally converted into control signals.
II-B2 Action Space $\mathcal{A}$
The action space $\mathcal{A}$ defines the set of possible actions that the platoon can take at any given time step. In this study, each autonomous vehicle in the platoon selects from a discrete set of high-level actions, including left turn, right turn, accelerate, decelerate, and cruise. These actions allow the vehicles to maneuver within their lanes and adjust their speeds to maintain a safe and efficient formation.
For a platoon consisting of $N$ autonomous vehicles, the joint action space $\mathcal{A}$ is the Cartesian product of the individual action spaces of the vehicles:
$\mathcal{A}=\mathcal{A}_{1}\times\mathcal{A}_{2}\times\cdots\times\mathcal{A}_{N}$ (5)
Each vehicle $i$ in the platoon selects an action $a_{i}\in\mathcal{A}_{i}$, and the combined action $a_{platoon}$ is executed as:
$a_{platoon}=(a_{1},a_{2},\ldots,a_{N})$ (6)
In this study, a quinary encoding scheme is employed to encode the joint action space, resulting in a 125-dimensional discrete action space through the Cartesian product of the individual actions. After selecting high-level actions, the platoon’s low-level controllers manage steering and throttle to execute the specific vehicle movements in the simulation environment, as illustrated in Figure 2.
II-B3 Transition Probabilities $\mathcal{P}$
HDVs introduce randomness into the whole system, making future states inherently uncertain. This uncertainty is modeled using a state transition function $\mathcal{P}(s_{t+1}|s_{t},a_{t})$, which represents the stochastic nature of the environment, capturing the likelihood of transitioning from the current state $s_{t}$ to the next state $s_{t+1}$ given the current action $a_{t}$:
$\mathcal{P}(s_{t+1}|s_{t},a_{t})=\Pr(s_{t+1}|s_{t},a_{t})$ (7)
The transition probabilities account for the stochasticity brought by the HDVs, ensuring that the decision-making process anticipates and responds to these uncertainties.
III Methodology
This section introduces a comprehensive framework for autonomous vehicle platoon. The DRL-based core decision-making algorithm for formation reconfiguration is first presented to enable adaptive and secure platoon adjustments, which is followed by a formation maintenance algorithm and an adaptive switching mechanism, ensuring stable and efficient decision-making across varying traffic conditions.
III-A Framework Overview
Figure 3: Overview of the TriCoD framework for self-organizing autonomous vehicle platooning, which combines a data-driven upper layer and a model-driven lower layer to manage platoon adaptively. The upper layer uses DRL with a twin-world safety verification model for safe, real-time decision-making, while the lower layer employs a LQR for precise control in predictable conditions, ensuring stable and efficient vehicle spacing. Together, these layers provide robust, adaptive platooning in complex traffic environments.
To address emergency and high-risk scenarios that traditional vehicle platooning technologies struggle to manage effectively, this study introduces TriCoD framework. As depicted in Figure 3, the framework seamlessly integrates two tightly coupled primary layers that collaboratively handle different aspects of platoon decision making:
  • Upper Layer: The DRL decision model follows the Centralized Training and Decentralized Execution (CTDE) paradigm, with a twin-world safety verification model that simulates decision outcomes in a parallel virtual environment. This verification ensures safe, effective decision-making before deployment and enhances the DRL model’s sampling efficiency. The upper layer works alongside the lower layer to adapt to complex scenarios in real-time.
  • Lower Layer: The Linear-Quadratic Regulator (LQR) provides precise control in predictable scenarios, minimizing headway for efficiency. This layer also collaborates with the DRL model by executing low-level control commands that align with the strategic decisions made by the upper layer.
A knowledge-based finite state machine dynamically integrates these layers, switching decision strategies based on real-time traffic conditions. In stable conditions, the LQR controller optimizes efficiency with reduced headway. In complex scenarios, the DRL model manages platoon splitting and reformation, coordinating with the LQR for adaptive responses. This hybrid approach combines data-driven learning with model-based control, ensuring cohesive, adaptive responses to both routine and emergency situations.
III-B Reward for Platooning
The reward function is central to shaping the behavior of autonomous vehicle platoons. Inspired by the social behavior in AVs[26], a reward system is designed to account for both individual vehicle performance and the collective dynamics of the platoon as a whole. This approach ensures that each vehicle operates safely and efficiently while maintaining coordination and cohesion within the platoon.
The total reward $R_{global}(s,a)$ is formulated as the sum of individual rewards and system-wide collaborative rewards:
$R_{global}(s,a)=\frac{1}{n}\sum_{i=1}^{n}R_{ind}^{i}+R_{sys}$ (8)
Both the individual vehicle perspective and the platoon-wide perspective are considered in the reward design. From the individual vehicle’s perspective, the reward function emphasizes safety and efficiency, focusing on collision avoidance, lane keeping, speed control, and smooth acceleration:
$R_{ind}(s,a)=\sum_{t\in\{C,F,A\}}w_{t}\times r_{t}$ (9)
where $r_{C}$ penalizes collisions, $r_{L}$ encourages lane discipline, $r_{F}$ rewards maintaining high speeds, and $r_{A}$ controls acceleration behavior. The weights $w_{C}$, $w_{L}$, $w_{F}$, and $w_{A}$ are tuned to reflect the importance of each factor, with a strong emphasis on collision avoidance through a higher value of $w_{C}$.
From the platoon-wide perspective, the reward function promotes the cooperation and cohesion of the entire platoon. The system-wide reward $R_{sys}(s,a)$ encourages vehicles to remain coordinated and maintain proper formation:
$R_{sys}(s,a)=\sum_{k\in\{M,D,H,S\}}w_{k}\times r_{k}$ (10)
In this expression, $r_{M}$ rewards staying in the same lane, $r_{D}$ encourages covering greater distances, $r_{H}$ ensures proper headway between vehicles, and $r_{S}$ aligns speed coordination within the platoon. The corresponding weights $w_{M}$, $w_{D}$, $w_{H}$, and $w_{S}$ are designed to ensure that each vehicle’s actions contribute to the collective success of the formation.
This reward design effectively balances the individual objectives of safety and efficiency with the collective need for coordination, ensuring that the platoon operates as a cohesive unit. The combined rewards guide both individual vehicles and the platoon as a whole toward achieving optimal performance in dynamic traffic conditions. Figure 4 illustrates the structure of the reward function, highlighting how it integrates both individual and collective factors.
Figure 4: Illustration of the designed headway maintaining reward, which considers both the single-vehicle rewards and the multi-vehicle rewards.
III-C Deep Reinforcement Learning Algorithm
Traditional DRL algorithms, such as Deep Q-Networks (DQN), have been applied in autonomous driving decision-making tasks. However, these algorithms primarily focus on single-agent scenarios and often struggle in multi-agent cooperative decision-making environments, such as vehicle platooning. This study addresses the challenge of reconstructing vehicle platoons in mixed traffic, aiming to develop a robust self-organizing decision model capable of handling highly dynamic and disruptive conditions through collaborative agent behavior.
To address this challenge, TRPO is selected for its ability to maintain stability by constraining policy updates within a trust region, thus preventing instability caused by drastic changes. This characteristic makes it well-suited for cooperative multi-vehicle decision-making, where stable and incremental improvements are critical.
A CTDE strategy is employed, as shown in Figure 5, where sensor data from all platooning vehicles are aggregated into a unified observation space. A shared policy $\pi_{\theta}$ then generates actions for all vehicles within the platoon. Following the execution of these actions by the low-level controllers, the policy $\pi_{\theta}$ is updated based on environmental feedback using experience replay and gradient-based optimization.
Figure 5: Illustration of the TRPO-based formation configuration algorithm schematic model.
In this self-organizing autonomous vehicle platooning system, all agents are treated equally and receive the same global reward. The objective of the DRL optimization is to maximize the global reward, defined as:
$\max{R_{g,t}=\sum_{i=1}^{N}r_{g,t}}$ (11)
Deep neural networks (DNNs) are employed to approximate the policy and value functions, enabling agents to learn how to make decisions within our environment to maximize cumulative rewards. This study employs the Actor-Critic architecture-based DRL neural network structure depicted in Figure 6.
Figure 6: Illustration of the Actor-Critic Network Structure with Security-Enhanced Action Probability, which processes inputs through full conneted layers in the Actor-Critic networks. The Actor generates action probabilities, which are refined by a security mask to ensure safe decisions, while the Critic evaluates actions to maximize cumulative rewards.
In this architecture, the policy loss and value function error loss are combined into a single loss function. Under the condition of shared parameters among agents, the total loss function can be expressed as a linear combination of the policy loss, state update loss, and entropy regularization term[27]:
$J\left(\theta_{i}\right)=J^{\pi_{\theta_{i}}}-\beta_{1}J^{V_{\phi_{i}}}+\beta_{2}H\left(\pi_{\theta_{i}}\left(s_{t}\right)\right)$ (12)
where $\beta_{1}$ and $\beta_{2}$ are the weight values for the value loss function and the entropy regularization term, respectively. The policy loss $J^{\pi_{\theta_{i}}}$ and the state update loss function $J^{V_{\phi_{i}}}$ are defined as:
$J^{\pi_{\theta_{i}}}=E_{\pi_{\theta_{i}}}\left[\log\pi_{\theta_{i}}\left(a_{i,t}|s_{i,t}\right)A_{i,t}^{\pi_{\theta_{i}}}\right]$ (13)
$J^{V_{\phi_{i}}}=\min_{\phi_{i}}E_{\mathcal{D}_{i}}\left[\left(r_{i,t}+\gamma V_{\phi_{i}}\left(s_{i,t+1}\right)-V_{\phi_{i}}\left(s_{i,t}\right)\right)\right]^{2}$ (14)
where $A_{i,t}^{\pi_{\theta_{i}}}$ is the advantage function, defined as:
$A_{i,t}^{\pi_{\theta_{i}}}=r_{i,t}+\gamma V^{\pi_{\phi_{i}}}(s_{i,t+1})-V^{\pi_{\phi_{i}}}(s_{i,t})$ (15)
where $V^{\pi_{\phi_{i}}}(s_{i,t})$ is the state value function. Additionally, the entropy regularization term encourages the agent to explore new directions and avoid potential local optima[28]. It is expressed as:
$H\left(\pi_{\theta_{i}}\left(s_{t}\right)\right)=E_{\pi_{\theta_{i}}}\left[-\log\left(\pi_{\theta_{i}}\left(s_{t}\right)\right)\right]$ (16)
This study updates the network parameters using mini-batch optimization sampling. The optimization objective is to minimize the variance of Equation 15.
III-D Twin World Security Projection
While DRL enables robust decision-making in mixed traffic with strong interferences, its ”black-box” nature can lead to unexpected outcomes in standard scenarios and may not fully comply with traffic regulations. Although more reward functions could enforce these constraints, they may incur significant training costs and risk model divergence. These issues are addressed in this study through the implementation of a comprehensive and extensible safety mask.
To ensure decision safety and sample efficiency, a safety-priority-based twin-world model is proposed. This model evaluates the safety of DRL decisions by first detecting potential collisions based on safety priorities, thereby reducing computational complexity. It then recursively computes $T_{n}$ steps in a twin world constructed from the current scenario, optimizing actions based on the platoon’s safety margin.
Given the focus of this study on decision-making in mixed traffic, predicting the driving behavior of HDVs is essential. The prevalent Intelligent Driver Model (IDM) [29] is utilized for simulating car-following behaviors, while the MOBIL model [30] is adopted for lane-changing decisions. The combination of these models minimizes collision risk among HDVs while closely mirroring human driving patterns.
The safety mask models each vehicle as a circle, with the center at the vehicle’s center and a radius equal to half its length. Future $T_{n}$ steps are simulated using motion models of HDVs and CAVs to assess collision risks. While this simulation is straightforward for individual vehicles, it becomes computationally challenging for platoons due to the exponentially growing action space. To address this, a safety-priority-based simulation model is proposed, which prioritizes vehicles with the smallest safety margins and conducts safety simulations in the twin world, ensuring efficient real-time computation and platoon cohesion.
Safety priority is determined based on headway time, with higher priority assigned to vehicles with smaller headway times relative to HDVs. The safety priority $p_{i}$ is defined as:
$p_{i}=-\log\frac{d_{headway}}{v_{t}}+\sigma_{i}$ (17)
where $d_{headway}$ is the headway distance, $v_{t}$ is the current vehicle speed, and $\sigma_{i}$ is a small random term to avoid identical safety priorities.
During the safety simulation, the vehicle with the highest safety priority undergoes assessment at each time step $t$. The twin world checks whether the vehicle’s decisions over the next $T_{n}$ time steps conflict with nearby vehicles. If no collision is detected (i.e., the minimum distance between vehicles exceeds the threshold), the DRL decision is deemed safe. Otherwise, the twin world traverses other possible decisions, selecting the action that maximizes safety margins:
$a_{t}^{\prime}=\arg\max_{a_{t}\in\mathcal{A}_{\mathrm{pro}}}\left(\min_{k\in T_{n}}d_{\mathrm{sm},k}\right)$ (18)
where $\mathcal{A}_{\mathrm{pro}}$ represents alternative actions output by the DRL, and $d_{\mathrm{sm}}$ refers to the safety margin, defined as the minimum headway distance in both current and target lanes during a left or right turn. The detailed procedure is illustrated in Algorithm 1.
Input: $vehicles$,$x_{i}$,$y_{i}$,$v_{x}^{i}$,$v_{y}^{i}$,$headway_{i}$
Output: $a_{i,i\in formation}$
1
2for $i=0$ to $N$ do
3       Compute safety priority: $priority_{i}=-\ln\left({headway_{i}}/{v_{i}}\right)$;
4       Sort vehicles by safety priority and store in list $\phi$;
5
6 end for
7for $j=0$ to $|\phi|$ do
8       Retrieve the vehicle with the highest safety priority: $\phi[0]$;
9       Identify neighboring vehicles and store them in list $N_{\phi[0]}$;
10       Predict the trajectory for the next $T_{n}$ simulation steps: $\xi_{v},v\in\phi[0]\cup N_{\phi[0]}$;
11       if trajectories overlap then
12             Calculate safety margin $d_{\text{sm},i}$;
13             if lane changed then
14                   $d_{\text{sm},i}=\min_{j}\left(p_{i}-p_{j},j\in\text{objective}\,v_{i}\right)$
15             end if
16            else
17                   $d_{\text{sm},i}=\text{headway}\,v_{i}$
18             end if
19            Replace action $a_{t}\leftarrow a_{t}^{\prime}$;
20             Update trajectory list $\xi_{\phi[0]}\leftarrow\xi_{\phi[0]}^{\prime}$;
21
22       end if
23      Replace action $a_{t}\leftarrow a_{t}^{\prime}$ Update $\phi[i]\leftarrow\phi[i+1],i=1,2,\cdots$
24 end for
return $a_{i,i\in formation}$
Algorithm 1 Twin World Security Projection
III-E Formation maintenance
Input: $vehicles$, $riskLevel$, $Participants$, $L_{safe}$
Output: $DecisionStrategy$
1
2for $i=0$ to $N$ do
3       Assess risk level and number of traffic participants within $L_{safe}$;
4       if ($riskLevel>=RoutineSafe$) and (Participants within $L_{safe}$) then
5             Switch to State $S_{1}$: LQR Control Model;
6             Activate LQR Control;
7             Set $DecisionStrategy=LQR$;
8
9      else
10             Switch to State $S_{2}$: Safety-Enhanced Data-Driven Strategy;
11             Set $DecisionStrategy=DataDriven$;
12
13       end if
14      Execute $DecisionStrategy$ based on the current state;
15       if $DecisionStrategy==LQR$ then
16             DecideWithLQR();
17
18       end if
19      else if $DecisionStrategy==DataDriven$ then
20             DecideWithDataDrivenStrategy();
21
22       end if
23
24 end for
return $DecisionStrategy$
Algorithm 2 Adaptive Switching Mechanism
The platoon reconfiguration model ensures safe, robust decision-making in complex, high-interference scenarios but struggles with maintaining close headway distances in routine conditions due to conservative safety strategies. To address this, an adaptive switching mechanism, the platoon maintenance model, is proposed, incorporating optimal control methods.
This model maintains close headway in routine scenarios while dynamically adjusting strategies based on real-time traffic conditions and interference levels, enhancing both safety in complex scenarios and efficiency in routine conditions.
To allocate decision strategies across scenarios, a Finite State Machine (FSM) $M(Q,D,\delta,q_{0},F)$ is used, where:
  1. 1.
    $Q=\{S_{1},S_{2}\}$ is the set of states;
  2. 2.
    $q\in Q$ represents the current state;
  3. 3.
    $\delta$ is the state transition function;
  4. 4.
    $F$ is the set of output states
  5. 5.
    $D$ denotes the risk level of the traffic scenario
This study considers a scenario as safe when the platoon is on the same lane and there are no other traffic participants within $L_{safe}$ ahead or behind the platoon. In this case, the FSM switches to state $S_{1}$, adopting the LQR control model to ensure smaller headway distances. In this study, string stability within the platoon[6] is not emphasized, as it has been thoroughly addressed in previous research. In all other scenarios, the FSM switches to state $S_{2}$, which employs a safety-enhanced data-driven strategy to balance the safety and efficiency of the platoon. By defining these states and transitions, the FSM provides a structured approach to adaptively manage the decision strategies of the platoon, ensuring optimal performance across varying traffic conditions. The detailed procedure is illustrated in Algorithm 2.
IV Simulation and Performance Evaluation
In this section, the software-in-the-loop (SIL) experiments are conducted to validate the effectiveness of the proposed method. Additionally, a hardware-in-the-loop (HIL) experimental platform is built to further verify the model’s safety and robustness and to assess its feasibility for direct deployment in real vehicles.
IV-A Software-in-the-loop experiment
IV-A1 Environment and simulation setting
Figure 7: Schematic representation of the model training reward function in multiple scenarios.
The simulation environment is constructed upon the Highway-Env simulator[31], a reinforcement learning experimental environment library based on an OpenAI Gym environment. As illustrated in Figure 2, all vehicles’ high-level policies are transformed into throttle and steering information through a low-level PID controller, thereby enabling vehicle movement within the simulation environment. Furthermore, our HDVs employ the IDM[29] car-following model and the MOBIL[30] lane-changing model provided by Highway-Env to achieve effective control in both longitudinal and lateral directions.
During the training process, a varying number of HDVs will randomly appear at randomly generated refresh points. The sampling frequency is set to $15$ Hz, meaning decisions and controls are made 15 times per second.
Three different random seeds are used to train the proposed model in various complex scenarios, with each session running for over 100,000 steps until convergence. Four algorithms—DQN [32], A2C [33], PPO [27], and RecurrentPPO [34]—are selected as experimental baselines. The training parameters for the DRL model are detailed in Table I.
TABLE I: Hyperparameter settings for reinforcement learning
Param Definition Value
$\eta$ Learning Rate $5\times{10}^{-4}$
$N_{steps}$ steps for each environment per update $256$
$N_{t}$ Total Training Steps $10^{5}$
$\beta_{1}$ Loss function weights 1
$\beta_{2}$ Loss function weights 0.01
$\lambda$ Minibatch size 128
$\gamma$ Discount factor 0.8
IV-A2 Performance evaluation
This section evaluates the performance of the proposed model within a simulated training environment, benchmarking it against advanced RL algorithms and conventional platooning strategies. Our findings indicate that the proposed platooning model consistently outperforms existing approaches across a variety of challenging scenarios.
The scenarios, depicted in Figure 1, are designed as edge cases to test the model’s safety and robustness under extreme conditions. The platoon vehicles travel at a speed of 28 m/s within a mixed traffic flow, where surrounding vehicles’ speed distribution follows a uniform range of 20 to 28 m/s. The traffic flow is characterized by oscillations and diverse driving styles, including aggressive, conservative, and neutral behaviors, creating a volatile environment prone to accidents.Additionally, malfunctioning vehicles may stall or brake unexpectedly, further increasing risk. The model is trained in this complex, high-risk scenario, with randomly generated traffic flows simulating real-world conditions.
Comparison with DRL baselines. Figure 7 illustrates the training performance of our model compared to baseline DRL algorithms across various scenarios. The training curves demonstrate that our model consistently outperforms the baseline algorithms, showing high adaptability and robustness essential for autonomous driving. This stability is attributed to our safety deduction model and innovative network structure, which enable our model to maintain reliable learning performance under challenging conditions and converge to a higher reward level.
Table II provides specific performance data. While our model may occasionally exhibit slightly lower speed and greater average headway distance compared to some baselines, it achieves significantly higher rewards, indicating that it prioritizes safety over raw efficiency. In contrast, the baseline algorithms reach their performance metrics at the expense of safety, which has been further verified in simulation experiments. Such trade-offs are unacceptable in transportation systems where safety is paramount.
In scenarios with high randomness and interference, such as traffic flow oscillations and traffic accidents, our model achieves significantly better convergence than other algorithms. This indicates an enhanced ability to learn optimal decisions under strong, dynamic disturbances, underscoring the robustness of our approach.
TABLE II: Performance metrics comparison between proposed method and benchmarks
Scenario Model Reward Avg. Speed Avg. HWD
Sce-1 A2C 18.93 22.23 10.00
DQN 9.47 25.18 20.06
PPO 22.10 23.83 13.85
RPPO 17.43 17.57 9.73
Ours 23.07 20.57 9.93
Sce-2 A2C 6.04 20.34 10.50
DQN 2.31 27.72 9.81
PPO 15.05 20.75 14.37
RPPO 4.10 25.70 17.10
Ours 31.66 20.59 11.33
Sce-3 A2C 4.87 20.18 10.00
DQN 11.83 20.71 12.18
PPO 58.21 20.13 10.00
RPPO 34.19 20.15 9.46
Ours 61.07 20.93 8.95
Comparison with Leading Platooning Algorithms. To evaluate the effectiveness of the TriCoD framework, a comparative analysis is conducted with several leading platooning algorithms. This analysis is based on four key performance indicators: average vehicle speed, average headway, collision rate, and pass rate. These metrics provide a comprehensive assessment of the model’s performance in terms of safety and efficiency. Notably, the pass rate metric is designed to consider scenarios where the convoy may need to halt to avoid hazards, thereby balancing risk avoidance with maintaining traffic flow.
In the experiments, the model is evaluated against advanced lane-changing algorithms, including SuPLC[35] and SiPLC[36], as well as a baseline DRL model across different scenarios: Human Inference, Traffic Accidents, and Traffic Flow Oscillation. As shown in Table III, our model consistently outperformed these benchmarks in terms of collision rate and pass rate, demonstrating a unique balance between safety and operational efficiency.
While some baseline models achieved slightly higher speeds or tighter headway distances, they do so at the expense of safety, as indicated by increased collision rates. In contrast, the TriCoD model maintained a zero-collision rate and a perfect pass rate of 100%, while still being able to maintain a relatively high pass rate and a smaller average headway. This balance illustrates our model’s capacity to prioritize safety without sacrificing operational effectiveness, especially in scenarios involving high-risk factors like traffic accidents and flow oscillations.
Overall, these results indicate that our model consistently achieves high safety and operational metrics across varied and challenging environments, demonstrating an effective balance that baseline models do not maintain.
TABLE III: Performance metrics comparison of multiple metrics
between the proposed method and three state-of-the-art benchmarks
Scenario Model Avg. Speed[m/s] Avg. HWD[m] Collision Rate Pass Rate
Human Inference SuPLC 22.58 9.68 1.00 0.00
SiPLC 22.58 9.68 1.00 0.00
DRL 20.46 10.00 0.00 1.00
Ours 20.57 9.93 0.00 1.00
Traffic Accidents SuPLC 21.96 10.18 0.52 0.29
SiPLC 20.81 10.15 0.54 0.21
DRL 21.13 14.10 0.19 0.91
Ours 20.59 11.33 0.00 1.00
Traffic Flow Oscillation SuPLC 20.28 10.13 0.03 1.00
SiPLC 19.68 10.15 0.03 0.97
DRL 20.93 10.08 0.09 0.91
Ours 20.93 8.95 0.00 1.00
IV-B Hardware-in-loop experiment
(a) HIL experimental platform framework
(b) HIL experiment platform physical diagram
Figure 8: Illustration of the HIL platform architecture for autonomous vehicle platooning, which integrates a CARLA simulation environment, RoadRunner for road scenes and traffic flow, and domin controllers for decision-making and control.
Figure 9: Illustration of the adaptive response to static obstruction in autonomous vehicle platooning, which decipts a scenario where the platoon encounters a multi-vehicle collision, necessitating coordinated lane changes to navigate the obstruction safely.
Figure 10: Illustration of the adaptive response to dynamic obstruction in autonomous vehicle platooning, which depicts a scenario where the platoon leader’s sudden deceleration triggers a staggered lane-change maneuver among following vehicles to prevent collisions and maintain traffic flow.
IV-B1 HIL experimental platform construction
In the HIL experiments, an advanced simulation platform is constructed by integrating a domain controller with Carla [37], a high-fidelity driving simulator, to validate the proposed model in a controlled yet realistic environment. The experimental highway scenario is created using Roadrunner, with varied background traffic flows generated by Highway-env.
TABLE IV: Specifications of the domain controller
Parameter Specification
CPU Cortex-A78AE – Arm®
CPU Frequency 2.2 GHz
Memory 64 GB 256-bit 204.8 GB/s
GPU Performance 275 TOPS
Storage 2 TB SSD
The platform consists of the domain controller, Carla simulation environment, Roadrunner, Highway-env, and PID controllers. The domain controller, whose parameters are shown in Table IV, processes real-time traffic and road information from Carla. Carla provides detailed road conditions and vehicle dynamics, which are translated into vehicle movements via PID controllers based on high-level action commands. Roadrunner generates complex highway scenarios, while Highway-env produces traffic with varied driving styles to test the algorithm’s robustness.
During experiments, the domain controller gathers real-time data from Carla, including vehicle positions, velocities, and traffic conditions.Based on this data, the proposed DRL algorithm makes high-level decisions, such as speed adjustments and lane changes. These decisions are sent back to Carla, where PID controllers convert them into specific throttle and steering commands. Various scenarios, including high-risk and normal conditions, are simulated to evaluate the algorithm’s performance based on metrics like safety, robustness, and efficiency. Figure 8 illustrates the platform’s overall architecture, highlighting its key components.
IV-B2 Case analysis
This section examines the performance of our model on the HIL simulation platform in real-time autonomous platooning scenarios. The results indicate that our model can effectively perform perception and decision-making, ensuring safe and robust platooning. Two typical scenarios with significant static and dynamic interferences are analyzed to illustrate the model’s strategies. Demonstration videos are available on our website  [^1] Hardware-in-the-Loop Experimental Validation Video Weblink.
In the first scenario, the platoon encounters a multi-vehicle collision blocking several lanes. Traditional strategies struggle with safe deceleration and coordination, but our model enables the lead vehicle to initiate a lane change, followed by others when safe. As shown in Figure 9, the lead vehicle moves at step 30, with all vehicles clearing the obstruction by step 220. The LQR controller stabilizes the platoon with reduced headway, illustrating flexibility in static obstructions. In the second scenario, the lead vehicle suddenly decelerates at 5 m/s² due to a breakdown. Traditional models risk collision due to short distances, but ours initiates a staggered lane change (see Figure 10). By step 70, all vehicles are safely repositioned, and the platoon reforms in the right lane by step 455, maintaining safety and flow.
These scenarios demonstrate the model’s capability to handle sudden obstacles by dynamically adjusting the platoon’s formation and leveraging available lane space to maintain safety and efficiency.
V Conclusion
In the evolving field of autonomous driving, ensuring safe and robust self-organizing decision-making for vehicle platooning, especially in mixed traffic, presents significant challenges. This paper introduces TriCoD, a Twin-World Safety-Enhanced Data-Model-Knowledge Triple-Driven Adaptive Autonomous Vehicle Platooning Cooperative Decision-Making Framework. By integrating reinforcement learning with model control techniques, TriCoD enables adaptive decision-making across diverse, complex scenarios, guided by traffic priors and optimal control strategies.
The proposed model significantly outperforms traditional platooning methods, such as CACC and CLCC, particularly in safety, efficiency, and robustness. Extensive simulations and HIL experiments validate its effectiveness in handling both static and dynamic disturbances.
Future work will focus on optimizing the model for scalability in larger platoons, refining its performance in real-world settings, and exploring the integration of additional sensory data to enhance situational awareness. These advancements will further advance the reliability and applicability of autonomous platooning systems in commercial deployments.